OpenStudy (anonymous):

which cosine function has maximum of .5 a minimum of -.5 and a period of (2pi/3)?

OpenStudy (anonymous):

@hermom1967 @e.mccormick @Hero

OpenStudy (anonymous):

@Frostbitten @ForeverNever

OpenStudy (anonymous):

@Frostbitten you there?

OpenStudy (anonymous):

Sorry, this is beyond me, I don't know the answer.

OpenStudy (anonymous):

It's ok

OpenStudy (anonymous):

@hero can you help me with this one?

OpenStudy (e.mccormick):

OK. Normally, cosine has a range of -10 to 1. What would change the range to -.5 TO .5?

OpenStudy (anonymous):

Um... would it be... ah I don't know. Can you help explain the procedure?

OpenStudy (anonymous):

*by explaining

OpenStudy (anonymous):

@e.mccormick ? Are you there?

OpenStudy (e.mccormick):

Yah. Just typing something up and my bos is asking me a question.

OpenStudy (anonymous):

oh wow, okay thank you. I didn't know you where at work :) I can wait

OpenStudy (anonymous):

Sorry, this is wwwwaaaaayyyyyyy beyond me. I can ask my friend though.

OpenStudy (e.mccormick):

For pretty much any function, trig or otherwise, the following is true: If the normal version is \(f(x)\) then transformations can be done with \(Af(Bx-h)+k\) A is the vertical stretching. B is the horizontal stretching h is the horizontal shift k is the vertical shift

OpenStudy (anonymous):

Okay, that makes sense. and okay @ForeverNever

OpenStudy (e.mccormick):

So, if you look at the unit circle, cosine has a range of -1 to 1. If A is the vertical stretching, and cosine is normally \(\cos (x)\) what would the A in \(A\cos (x)\) be to change the range to -.5 to .5?

OpenStudy (anonymous):

Wouldn't it be .5?

OpenStudy (e.mccormick):

Yes! So that is part of the answer: \(.5 \cos(x)\) is where we are at and that is half the work. The period is the horizontal domain. The normal period is \(2\pi\). Now, changes in period are inverted. That means that to do a period that is twice as long, you multiply by \(\frac{1}{2}\). To do a period that is half as long, you multiply by 2. So, what would change \(2\pi\) to \(\dfrac{2\pi}{3}\) and then what is the inverse of that?

OpenStudy (anonymous):

Sorry, he doesn't know. all he said was his brain exploded when he read it

OpenStudy (anonymous):

Umm, dividing by three and the inverse is... I don't know, isn't it just the reverse? It's okay @ForeverNever

OpenStudy (e.mccormick):

If I have \(\dfrac{5}{a}\) then the inverse is \(\dfrac{a}{5}\). So you flip the fraction.

OpenStudy (anonymous):

Ok so yeah, it's just the reverse. So 3/2pi?

OpenStudy (e.mccormick):

Well, you only want what would change it. So what would change it is \(\dfrac{1}{3}\) and the inverse is \(\dfrac{3}{1}\) or just 3.

OpenStudy (anonymous):

Okay. Yeah I get that...

OpenStudy (e.mccormick):

So, where would the 3 need to go. Just look at my list of rules and what I said it is changing and see what you think.

OpenStudy (anonymous):

Well, If we multiply 2*3.14, we get around 6. So 3/6=1/3?

OpenStudy (e.mccormick):

Well, I mean where in the \(.5cos(x)\) does the 3 need to go? That is all that is left. You already changed the range, so all you need it to finish changing the period and it is done.

OpenStudy (anonymous):

Well the three goes in the x right?

OpenStudy (e.mccormick):

By in the x you mean in place of or next to?

OpenStudy (anonymous):

In the place of it.

OpenStudy (e.mccormick):

No, the x still needs to be there or it is no longer a function, just an equation.

OpenStudy (e.mccormick):

AH, if you want to sue \(\theta\) instead of x, then yes.

OpenStudy (e.mccormick):

Use... not sue. Ignore the dyslexia. You just need a variable, and \(\theta\) works.

OpenStudy (anonymous):

Okay, that makes sense. Thank you for helping me :)

OpenStudy (e.mccormick):

No problem. Have fun! Oh, and those rules, with a little extra info on what -A and -B do, are valid for most anything. So the things that shift a parabola or a circle or just about anything will be the same or very similar.

OpenStudy (anonymous):

Okay, I will keep that in mind :)

OpenStudy (e.mccormick):

-A flips over the x axis. -B flips over the y axis. That is the only other thing to know. Here are some examples that are from non-trig functions and you can see the numbers in the same places doing the same sort of changes: http://tutorial.math.lamar.edu/Classes/Alg/Transformations.aspx

OpenStudy (anonymous):

Okay I will check the link out :) Thank you!

OpenStudy (e.mccormick):

=) Welcome!